Find the mass of the ammonia in the 100 ft^3 3 tank having a pressure of 70 psi at 120
1
Expert's answer
2021-10-10T15:55:06-0400

"V=100~ft^3=2.83~m^3,"

"p=70~psi=4.83\cdot 10^5~Pa,"

"T=70\u00b0F=294~K,"

"M=17\cdot10^{-3}~\frac{kg}{mol},"

"pV=\frac mM RT,\implies"

"m=\frac{pVM}{RT}=9.5~kg."